Please help me determine the slope and the y-intercept; y= -18x-772
Found 2 solutions by solver91311, drj:
Answer by solver91311(24713)



Since your equation is already in slope-intercept form, , you can determine the slope and -intercept by inspection. The number that is in the place of the in is the slope, and the number that is in the place of the in is the -intercept.


John


Answer by drj(1380)

Please help me determine the slope and the y-intercept; y= -18x-772

Step 1. The slope-intercept form of a linear equation is given as y=mx+b where m is the slope and b is the y-intercept when x=0 or at point (0,b). The given equation y=-18x-772 is in slope-intercept form.

Step 2. ANSWER: Comparing y=-18x-772 with y=mx+b, then the slope m=-18 and b=-772.
I hope the above steps and explanation were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J
http://www.FreedomUniversity.TV